# StyleMaster: Towards Flexible Stylized Image Generation with Diffusion Models

### v1

@@ 4.2 Dynamic attention adaptation
The extracted style embedding $z_{s}$ from $\mathcal{I}_{s}$ as described above can then be used to adapt the pretrained SD to guide the denoising process based on style information. Thanks to the design of self-attention and cross-attention mechanism in diffusion UNet, such adaptation can be simply instantiated as an extra cross-attention module that is parallel to the original prompt-based cross-attention, as adopted in StyleAdapter. However, we empirically find that such a method is suboptimal in a large amount of cases, leading to severe semantic inconsistency. A straightforward solution is cut down the number of extra attention modules so that only upsample layers in diffusion UNet are adapted. In this way, the text prompt can dominate the cross-attention in half of the UNet, consequently resulting in better semantics in the generated images. However, this can decrease the capacity of adapters, leading to less preferable stylization. To this end, we propose to adopt a dynamic adaptation strategy which is applied to both self-attention and cross-attention (Fig. 3(b)).
Dynamic self-attention adapter. As discussed in previous works [8], the projected value tensor $W_{V}^{l}\cdot z^{l}$ in the self-attentions contributes to the texture of generated images. Therefore we introduce a dynamic self-attention adapter module based on adaIN. Formally, for the $l$-th self-attention layer, we project $z_{s}$ with a linear layer and adjust $W_{V}^{l}\cdot z^{l}$ according to statistics of $z_{s}$:
$\hat{\mathbf{V}}^{l}=\mu(f_{proj-SA}^{l}(z_{s}))+\frac{\sigma(f_{proj-SA}^{l}(z_{s}))}{\sigma(W_{V}^{l}\cdot z^{l})}*(W_{V}^{l}\cdot z^{l}-\mu(W_{V}^{l}\cdot z^{l}))$ (3)
where $f_{proj-SA}^{l}$ denotes dynamic projection layer, $\mu,\sigma$ denote mean and standard deviation. By rescaling $\mathbf{V}^{l}$, the information contained in $z_{s}$ can be directly embedded into the image feature without destroying the structure and semantic meaning of the generated image.
Dynamic cross-attention adapter. To adapt the cross-attention layers, we follow the idea of StyleAdapter to adopt the dual-path cross-attention mechanism. Basically, for the $l$-th cross-attention layer, besides the original cross-attention performed between text embedding $z_{text}$ and image embedding $z^{l}$ as in Sec. 3, an extra style-aware cross-attention is added as $\tilde{z}^{l}=Attention(\hat{W}_{Q_{t}}^{l}\cdot z^{l},\hat{W}_{K_{s}}^{l}\cdot z_{s},\hat{W}_{V_{s}}^{l}\cdot z_{s})$ with additional learnable parameters $\hat{W}_{K_{s}}^{l},\hat{W}_{V_{s}}^{l}$. Then $\hat{z}^{l}+\lambda\tilde{z}^{l}$ is fed into the following feed-forward networks, where $\lambda$ is learnable coefficient.
@@
where $f_{proj-CA}$ denotes a linear projection layer, $\cdot\|_{c}\cdot$ denotes concatenation along the channel dimension. Then two weight generators instantiated as linear layers are applied to $z_{s}^{l}$, resulted in two dynamic weights $W_{K_{d}}^{l},W_{V_{d}}^{l}$ with dimension $d*d^{l}$. These two features are reshaped into linear layer weights and used to transform $z_{s}$. After that the style-aware cross-attention is modified as
$\tilde{z}^{l}=Attention(\hat{W}_{Q_{t}}^{l}\cdot z^{l},(\hat{W}_{K_{s}}^{l}+W_{K_{d}}^{l})\cdot z_{s},(\hat{W}_{V_{s}}^{l}+W_{V_{d}})\cdot z_{s})$ (5)
In this way, the key and value projections are partially dependent on $z_{s}$, resulting in a more complex transformation of $z_{s}$ and leading to better capacity. To make this module parameter-efficient, we adopt a grouping strategy, i.e.channels of $z_{s}$ in each group share the same dynamic weight to produce $W_{K_{d}}^{l}$ and $W_{V_{d}}^{l}$, hence lighter weight generators can be used to generate dynamic weights.
@@ 4.3 Training objectives
where $\mathcal{L}_{noise}$ is the noise prediction loss as in Eq. 1. $\mathcal{L}_{disen}$ denotes a semantic disentangle loss applied to style embedding $z_{s}$. To ensure that the proposed style embedding model can get rid of the semantic information contained in $\mathbf{I}_{style}$ when producing $z_{s}$, $\mathcal{L}_{disen}$ is designed by enlarging the similarity between $z_{s}$ and $z_{CLIP}$ while decreasing the similarity between $z_{s}$ and text embedding $z_{text}$. As discussed in Sec. 4.1, the text embedding represents more abstract semantic information than the image embedding. Therefore this loss term can help the model avoid the possibility of semantic leakage, thus leading to better style embedding. On the other hand, to enhance the style consistency, we propose to regulate the gram matrix of $\hat{x}^{0}$, which is the noisy estimation from $z^{t}$ and can be calculated as
@@
Specifically, we apply several rigid transformations such as random rotation and cropping to $\mathcal{I}_{S}$ to get a new image $\mathbf{I}_{pos}$ that has the same style as $\hat{x}_{0}$. Then elastic transformation and color jitter are also applied to $\mathcal{I}_{S}$. The resulted $\mathbf{I}_{neg}$, while sharing similar semantic object to $\mathbf{I}_{inp}$, barely inherits the style from it. Then the objective is a triplet loss which can be written as
@@
where $\mathcal{G}$ denotes the gram matrix of features. By optimizing this loss term, the model is encouraged to learn more detailed style information, thus leading to better results.
@@ 5.1 Implementation d etail
Experiment setting. Our experiments cover both one-shot and multi-shot settings. To make the evaluation more challenging, the shot number, i.e.number of reference images, varies from 2 to 5 among different styles in the multi-shot setting. We use all 20 styles for multi-shot experiments and 10 of them for one-shot experiments.
Training details. We adopt AdamW as optimizer with 1e-5 learning rate. Our model is trained for 200,000 iterations on 8 V100s with 8 batch size on each gpu.
Competitor. We include extensive methods as our competitor. For 1-shot experiment, MicroAST [20], StyleAlign [8], StyTR2 [4], and StyleAdapter [21] are adopted. For multi-shot experiment, InST [26], LoRA [10], Textual Inversion (TI) [5] and StyleAdapter are adopted.
5.2 Quantitative results
Table 1: Quantitative results for multi-shot setting. For all metrics, the larger score denotes the better model.
1-shot Methods Text Sim $\uparrow$ StyleSim $\uparrow$
MicroAST [20] 0.299 0.529
StyleAlign [8] 0.276 0.645
StyTR2 [4] 0.298 0.541
@@ 5.2 Quantitative results
Table 2: Quantitative results for multi-shot setting. For all metrics, the larger score denotes the better model.
multi-shot Methods Text Sim $\uparrow$ StyleSim $\uparrow$
InST [26] 0.196 0.692
@@
Objective quantitative results. For quantitative evaluation we adopt CLIP to calculate the style similarity between generated images and target style reference images, and the semantic similarity between generated images and target text prompts. The results are presented in Tab. 2 and Tab. 2 for 1-shot and multi-shot respectively. Specifically, MicroAST and StyTR2 share similar text similarity to ours one-shot experiments. This is because we use pretrained SD to generate base images for them, thus the basic semantic meaning is contained in the image. However the gap in terms of style similarity is much more marginal. Our method outperforms the best baseline StyleAdapter by 0.04. The results for multi-shot setting is consistent with one-shot, in which our method is significantly better then the competitors, thus showing the superiority of the proposed method.
Table 3: User Study for multi-shot setting. For all metrics, the larger score denotes the better model.
1-shot Methods Text Sim StyleSim
MicroAST 3.18 2.84
StyleAlign 2.97 2.83
StyTR2 3.85 3.25
StyleAdapter 3.16 3.30
Ours 3.20 3.80
Table 4: User Study for multi-shot setting. For all metrics, the larger score denotes the better model.
multi-shot Methods Text Sim StyleSim
InST 2.11 2.68
LoRA 2.84 3.34
TI 2.95 3.03
StyleAdapter 3.12 3.78
Ours 3.14 3.86
Subjective quantitative results. Apart from the quantitative results in the main paper, we randomly sample 5 images for each method and each style and conduct a user study. During the study, the participant are required to score 1-5 for each image by considering both semantic fidelity and style consistency. Then all scores are averaged. The results are shown in Tab. 4 and Tab. 4, which are generally consistent with the objective results except that the text similarity of StyTR2 is extremely high with human scores. We think this may be attributed to the fact that our method still lacks behind the SD model used to generated base images for StyTR2 in terms of semantic fidelity. On the other hand, while StyTR2 receives good semantic fidelity, its style consistency is poor. Our method, on the contrary, can find a good balance for the trade off between semantic and style.
@@ 5.3 Qualitative results
Figure 4: One-shot qualitative comparison. Styles from top to bottom: Cezanne, flat cartoon, expressionism, ink. Prompts from left to right: A bird in a word; A boy wearing glasses, he is reading a thick book; A cherry blossom. For detailed reference images please refer to the appendix.
Figure 5: Multi-shot qualitative comparison. Styles from top to bottom: pencil, watercolor, Monet, impasto. Prompts from left to right: A robot; A modern house with a pool; A lake with calm water and reflections. For detailed reference images please refer to the appendix.
@@
We present several uncurated results for both settings in Fig. 4 and Fig. 5. For one-shot experiment, since we use pretrained SD to first generate the content images, the style transfer based methods such as StyTR2 and MicroAST can generally enjoy reasonable semantic consistency. However, such methods can only inherit the basic color information from reference images rather than the detailed style information such as shape, texture and layout, thus making them less preferable. For example, they fail to present the cubism and the curves in the second and third row. This is because these methods rely on simple representation to transfer the style-related knowledge from reference images, which leads to the problem of under-stylization. StyleAlign, on the other hand, can stylize the images better. For example the images generated by StyleAlign in the fourth row share similar style patterns with the reference image. However, we find that StyleAlign suffers from inconsistent semantics problem, which leads to inconsistent semantic meaning with the text prompt. Moreover, images generated by StyleAlign seems messy and disordered, which may be attributed to the uncontrolled shared self-attention between content and style images. The performance of StyleAdapter is much better than other competitors, while it is hard for this method to understand complex style patterns, leading to undesirable results when it comes to ink painting (fourth row). Compared with these methods, our method can learn appropriate style information from reference images, e.g. the scattered color patches in the first and fourth row, and simultaneously keep the images faithful to the prompts, thus making the best of both worlds.
@@
The multi-shot setting which is more challenging shows similar results. InST can hardly replicate the style. LoRA and TI suffer from limited style information. StyleAdapter, while utilizing a specifically-designed pipeline, shows a tendency to confuse the given styles with the photographic prior knowledge from pretrained SD. Such phenomenon is most obvious in the first row of Fig. 5, where pencil drawings are provided as reference images, but StyleAdapter generates grayscale photos. Our method, thanks to the proposed multi-source style embedder which can extract more detailed style information and the dynamic attention adaptation, can generally generate different kinds of styles with high image quality and semantic fidelity.
@@
5.4 Ablation study
To further verify the efficacy of our contributions, we conduct several ablation studies on multi-shot setting. More quantitative and qualitative ablation studies are provided in the appendix.
Design of style embedding module. We consider two variants together with the full model for the style embedding module: not using the VGG gram matrix to regulate the attention layers (w/o VGGToken), and not engaging the reference image captions in the extraction (w/o NegEmb). The results are shown in the left three columns of Fig. 6. The style of images generated by model without VGGToken is generally less mimic. Meanwhile, the model without NegEmb not only has worse style (first and second row) but also suffers from wrong semantic meaning (third row).
Design of attention adapter. We illustrate the role of different parts in the proposed dynamic attention adapter in the middle three columns of Fig. 6. When the model adopts attention adapter for all UNet attention layers instead of only the upsample ones, the images generally have problem of wrong semantic meaning. For example, the clock is missing in the first row, and the cloth color is wrong in the third row. Also, it is obvious that when only using the same cross-attention adapter as in StyleAdapter, the generated images show inconsistent and undesirable styles, which can be attributed to the limited capacity to such strategy.
Effectiveness of different objective functions. In the right three columns of Fig. 6 we inspect the efficacy of two objectives introduced in Sec. 4.3. The results directly support our claim that the gram consistency loss can enhance the style in generated images and the semantic disentangle loss can make the model tell apart semantic and style information from reference images, thus better handling contents in the text prompts.

### v2

@@ 4.2 Dynamic Attention Adaptation
The extracted style embedding $z_{s}$ from $\mathcal{I}_{s}$ as described above can then be used to adapt the pretrained SD to guide the denoising process based on style information. Thanks to the design of self-attention and cross-attention mechanism in diffusion UNet, such adaptation can be simply instantiated as an extra cross-attention module that is parallel to the original prompt-based cross-attention, as adopted in StyleAdapter. However, we empirically find that such a method is suboptimal in a large amount of cases, leading to severe semantic inconsistency. A straightforward solution is cut down the number of extra attention modules so that only upsample layers in diffusion UNet are adapted, as in DEADiff and InstantStyle. In this way, the text prompt can dominate the cross-attention in half of the UNet, consequently resulting in better semantics in the generated images. However, this can decrease the capacity of adapters, leading to less preferable stylization. To this end, we propose adopting a Dynamic Attention Adaptation (DAA) strategy which is applied to both self-attention and cross-attention (Fig. 4).
Dynamic self-attention adapter. As discussed in previous works [9], the projected value tensor $W_{V}^{l}z^{l}$ in the self-attentions contributes to the texture of generated images. Therefore we introduce a dynamic self-attention adapter module based on adaIN. Formally, for the $l$-th self-attention layer, we project $z_{s}$ with a linear layer and adjust $W_{V}^{l}z^{l}$ according to statistics of $z_{s}$:
$\hat{\mathbf{V}}^{l}=\mu(f_{SA}^{l}(z_{s}))+\frac{\sigma(f_{SA}^{l}(z_{s}))}{\sigma(W_{V}^{l}z^{l})}(W_{V}^{l}z^{l}-\mu(W_{V}^{l}z^{l}))$ (3)
where $f_{SA}^{l}$ denotes dynamic projection layer, $\mu,\sigma$ denote mean and standard deviation. By rescaling $\mathbf{V}^{l}$, the information contained in $z_{s}$ can be directly embedded into the image feature without destroying the structure and semantic meaning of the generated image.
Dynamic cross-attention adapter. To adapt the cross-attention layers, we follow the idea of StyleAdapter to adopt the dual-path cross-attention mechanism. Basically, for the $l$-th cross-attention layer, besides the original cross-attention performed between text embedding $z_{text}$ and image embedding $z^{l}$ as in Sec. 3, an extra style-aware cross-attention is added as $\tilde{z}^{l}=Attn(\hat{W}_{Q_{t}}^{l}z^{l},\hat{W}_{K_{s}}^{l}z_{s},\hat{W}_{V_{s}}^{l}z_{s})$ with additional learnable parameters $\hat{W}_{K_{s}}^{l},\hat{W}_{V_{s}}^{l}$. Then $\hat{z}^{l}+\lambda\tilde{z}^{l}$ is fed into the following feed-forward networks, where $\lambda$ is learnable coefficient.
@@
where $f_{proj-CA}$ denotes a linear projection layer, $\|_{c}$ denotes concatenation along the channel dimension. Then two weight generators instantiated as linear layers are applied to $z_{s}^{l}$, resulted in two dynamic weights $W_{K_{d}}^{l},W_{V_{d}}^{l}$ with dimension $d*d^{l}$. These two features are reshaped into linear layer weights and used to transform $z_{s}$. After that the style-aware cross-attention is modified as
$\tilde{z}^{l}=Attn(\hat{W}_{Q_{t}}^{l}z^{l},(\hat{W}_{K_{s}}^{l}+W_{K_{d}}^{l})z_{s},(\hat{W}_{V_{s}}^{l}+W_{V_{d}}^{l})z_{s})$ (5)
In this way, the key and value projections are partially dependent on $z_{s}$, resulting in a more complex transformation of $z_{s}$ and leading to better capacity. To make this module parameter-efficient, we adopt a grouping strategy, i.e., channels of $z_{s}$ in each group share the same dynamic weight to produce $W_{K_{d}}^{l}$ and $W_{V_{d}}^{l}$, hence lighter weight generators can be used to generate dynamic weights.
@@ 4.3 Training Objectives
where $\mathcal{L}_{noise}$ is the noise prediction loss as in Eq. 1. $\mathcal{L}_{disen}$ denotes a semantic disentangle loss applied to style embedding $z_{s}$. To ensure that the proposed style embedding model can get rid of the semantic information contained in $\mathbf{I}_{style}$ when producing $z_{s}$, $\mathcal{L}_{disen}$ is designed by enlarging the similarity between $z_{s}$ and $z_{CLIP}$ while decreasing the similarity between $z_{s}$ and text embedding $z_{text}$. Formally, $L_{disen}=sim(z_{cap},z_{s})-\delta sim(z_{CLIP},z_{s})$, where $\delta$ is hyper-parameter set as 0.1, $sim$ denotes cosine similarity. As discussed in Sec. 4.1, the text embedding represents more abstract semantic information than the image embedding. Therefore this loss term can help the model avoid the possibility of semantic leakage, thus leading to better style embedding. On the other hand, to enhance the style consistency, we propose to regulate the Gram matrix of $\hat{x}^{0}$, which is the noisy estimation from $z^{t}$ and can be calculated as
@@
Specifically, we apply several rigid transformations such as random rotation and cropping to $\mathcal{I}_{S}$ to get a new image $\mathbf{I}_{pos}$ that has the same style as $\hat{x}_{0}$. Then elastic transformation and color jitter are also applied as style destroy method to $\mathcal{I}_{S}$. The resulted $\mathbf{I}_{neg}$, while sharing similar semantic object to $\mathbf{I}_{inp}$, barely inherits the style from it. Then the objective is a triplet loss which can be written as
@@
where $\mathcal{G}$ denotes the Gram matrix of features. By optimizing this loss term, the model is encouraged to learn more detailed style information, thus leading to better results. In total, during training, only the style embedder and the added adapters are trained with objectives as in Eq. 6. Those used backbones such as VGG, CLIP and original parameters from SD are not trained. During inference, we directly use the trained models to generate stylized images without any test-time optimization.
@@ 5.1 Implementation Detail
Experiment setting. Our experiments cover both one-shot and multi-shot settings. To make the evaluation more challenging, the number of reference images varies from 2 to 5 among different styles in multi-shot setting. We use all 20 styles for multi-shot experiments and 10 of them for one-shot experiments. Our proposed method does not need test-time optimization for both settings. In multi-shot experiment, the style descriptors produced by the MSD for all reference images are concatenated as the input condition.
Training details. We adopt AdamW as optimizer with 1e-5 learning rate. Our model is trained for 200,000 iterations on 8 V100s with 8 batch size on each gpu, which takes about 3 days. Our code will be release.
Competitor. We include extensive methods as our competitor. For 1-shot experiment, MicroAST (MAST) [26], StyleAdapter (SAda) [25], StyleDrop (SDrop) [21], InstantStyle (InsStyle) [24], DEADiff [17] and StyleID [5] are adopted. For multi-shot experiment, InST [30], LoRA [11], Textual Inversion (TI) [6], StyleDrop [21], StyleAdapter and InstantStyle are adopted. Among all methods, MicroAST, InST, LoRA, and TI require test-time optimization, and the others are finetune-free methods.
Table 1: Quantitative results for one-shot setting.
Backbone 1-shot Methods Text Sim $\uparrow$ Style Sim $\uparrow$
SD15 MicroAST [26] 0.299 0.529
StyleDrop [21] 0.290 0.583
InstantStyle [24] 0.167 0.840
@@
SDXL StyleAlign [9] 0.276 0.645
InstantStyle [24] 0.295 0.652
Ours 0.311 0.696
Table 2: Quantitative results for multi-shot setting.
Backbone multi-shot Methods Text Sim $\uparrow$ Style Sim $\uparrow$
SD15 InST [30] 0.196 0.692
@@
SDXL InstantStyle [24] 0.291 0.645
Ours 0.293 0.667
Figure 5: One-shot qualitative comparison with SD1.5 as backbone. For comparison with SDXL as backbone, please refer to the supplementary material. Zoom in for more details.
Figure 6: Multi-shot qualitative comparison with SD1.5 as backbone. For detailed reference images and comparison with SDXL as backbone, please refer to the supplementary material. Zoom in for more details.
5.2 Quantitative Results
For quantitative evaluation we adopt CLIP to calculate the style similarity between generated images and target style reference images, and the semantic similarity between generated images and target text prompts. The results are presented in Tab. 2 and Tab. 2 for 1-shot and multi-shot respectively. Note that for SD1.5, InstantStyle receives incredibly high style similarity, which is not attributed to its strong performance. In fact, as we will show in the qualitative results, InstantStyle-SD1.5 totally leaks the content in the reference images into the generated images, thus leading to extremely poor text similarity. Moreover, MicroAST shares similar text similarity to ours one-shot experiments. This is because we use pretrained SD to generate base images for them, thus the basic semantic meaning is contained in the image. However the gap in terms of style similarity is much more marginal. As for SDXL, our method performs generally better than other competitors. The results for multi-shot setting are consistent with one-shot, thus showing the superiority of the proposed method.
@@ 5.3 Qualitative Results
We present several uncurated results with SD1.5 as backbone for both settings in Fig. 5 and Fig. 6 respectively. First of all, as mentioned above, we find that InstantStyle suffers from problem of content leakage, resulting in meaningless generation. For one-shot experiment, since we use pretrained SD to first generate the content images, the style transfer based methods such as MicroAST can generally enjoy reasonable semantic consistency. However, such methods can only inherit the basic color information from reference images rather than the detailed style information such as shape, texture and layout, thus making them less preferable. For example, they fail to present the textures and curves. This is because these methods rely on simple representation to transfer the style-related knowledge from reference images, which leads to the problem of under-stylization. DEADiff, on the other hand, can stylize the images better. However, we find that the generated images of DEADiff, while being visually approvable, cannot follow the target style as shown in the reference images. Moreover, images generated by DEADiff seems to rely more on its learned prior knowledge rather than the input condition. The performance of StyleAdapter is generally reasonable, while it is hard for this method to understand complex style patterns, leading to undesirable results when it comes to ink painting. Compared with these methods, our method can learn appropriate style information from reference images, e.g., the scattered color patches in the first and fourth row, and simultaneously keep the images faithful to the prompts, thus making the best of both worlds.
@@
The multi-shot setting which is more challenging shows similar results. InST can hardly replicate the style. LoRA and TI suffer from limited style information. StyleAdapter, while utilizing a specifically-designed pipeline, shows a tendency to confuse the given styles with the photographic prior knowledge from pretrained SD. Such phenomenon is most obvious in the first row of Fig. 6, where pencil drawings are provided as reference images, but StyleAdapter generates grayscale photos. Our method, thanks to the proposed MSD which can extract more detailed style information and the dynamic attention adaptation, can generally generate different kinds of styles with high image quality and semantic fidelity.
Table 3: Ablation study for both one-shot and multi-shot settings.
Methods One-shot Multi-shot
Text Sim Style Sim Text Sim Style Sim
w/o Gram 0.288 0.692 0.288 0.698
w/o NegEmb 0.285 0.689 0.281 0.696
w/o OnlyUP 0.286 0.697 0.286 0.702
w/o DA 0.301 0.611 0.299 0.643
w/o $\mathcal{L}_{style}$ 0.280 0.694 0.290 0.705
w/o $\mathcal{L}_{disen}$ 0.282 0.695 0.286 0.694
Ours 0.299 0.708 0.291 0.719
5.4 Ablation Study
To further verify the efficacy of our contributions, we conduct several ablation studies on multi-shot setting. The quantitative results are shown in Tab. 3. More qualitative ablation studies are provided in the supplementary material.
@@ 5.4 Ablation Study
Design of style embedding module. We consider two variants together with the full model for the style embedding module: not using the Gram-based descriptor to regulate the attention layers (w/o Gram), and not engaging the semantic descriptor (w/o NegEmb). The results are shown in Fig. 7. The style of images generated by model without Gram is generally less mimic. Meanwhile, the model without NegEmb not only has worse style but also suffers from mistaken semantic meaning.
Figure 8: Images generated by different model variants.
Design of attention adapter. We illustrate the role of different parts in the proposed dynamic attention adapter in Fig. 8. When the model adopts attention adapter for all UNet attention layers instead of only the upsample ones, the images generally have problem of mistaken semantic meaning. For example, the clock is missing in the first row, and the cloth color is mistaken in the second row. Also, it is obvious that when only using the same cross-attention adapter as in StyleAdapter, the generated images show inconsistent and undesirable styles, which can be attributed to the limited capacity to such strategy. Interestingly, we find in Tab. 3 that when not using dynamic adapter, the model has a very different tendency compared with the full model, with much better text prompt similarity but much worse style similarity. This is reasonable since the proposed dynamic adapter greatly strengthens the impact of style embedding during both self-attention and cross-attention. In this way, the cross-attention with text prompts is weakened is disguise. In general, adopting dynamic adapter can make a good balance between text prompt fidelity and style fidelity.
Figure 9: Images generated by different model variants.
Effectiveness of different objective functions. In Fig. 9 we inspect the efficacy of two objectives introduced in Sec. 4.3. The results directly support our claim that the gram consistency loss can enhance the style in generated images and the semantic disentangle loss can make the model tell apart semantic and style information from reference images, thus better handling contents in the text prompts. Note that in one-shot setting, using $\mathcal{L}_{style}$ leads to larger improvement in terms of text similarity compared with using $\mathcal{L}_{disen}$. We think this may be because $\mathcal{L}_{style}$ may also help model to extract better style descriptor so that the generated images share more similar styles with the reference ones, thus leading to higher text similarity in this setting.